### v1

# Effective grand-canonical description of condensation in negative-temperature regimes

Abstract
The observation of negative-temperature states in the localized phase of the the Discrete Nonlinear Schrödinger (DNLS) equation has challenged statistical mechanics for a long time. For isolated systems, they can emerge as stationary extended states through a large-deviation mechanism occurring for finite sizes, while they are formally unstable in grand-canonical setups, being associated to an unlimited growth of the condensed fraction. Here, we show that negative-temperature states in open setups are metastable but their lifetime $\tau$ is exponentially long with the temperature, $\tau\approx\exp(\lambda|T|)$ (for $T<0$ and $\lambda>0$). More precisely, we find that condensation on a given site (i.e. emergence of a tall discrete breather) can advance only once a critical mass has been accumulated therein. This result has been obtained by combining the development of an effective grand-canonical formalism with the implementation of suitable heat baths. A general expression for $\lambda$ is obtained in the case of a simplified stochastic model of non-interacting particles. In the DNLS model, the presence of an adiabatic invariant, makes $\lambda$ even larger because of the resulting freezing of the breather dynamics. This mechanism, based on the existence of two conservation laws, provides a new perspective over the statistical description of condensation processes.
Statistical systems exhibiting negative (absolute) temperatures (NTs) often display awkward physical properties, but a careful analysis shows that their behavior does not contradict any fundamental law of thermodynamics Baldovin et al. (2021). NTs typically emerge when the total energy is somehow limited from above Ramsey (1956), but they can occur also in systems with unbounded energies, provided that the first moment of the density of states in energy remains finite Machlup (1975). Lasers are a clear example of the first kind: maximal energy is attained only by setting all the atoms in the excited state. Vortices in bounded fluids are an example of the second kind Onsager (1949); Gauthier et al. (2019); Johnstone et al. (2019). Other recent observations of NT states include cold atoms in optical lattices Braun et al. (2013) and multimode optical fibers Baudin et al. (2023); Marques Muniz et al. (2023).
Energy limitation can also arise indirectly in the presence of additional conservation laws. The discrete nonlinear Schrödinger (DNLS) equation, the subject of this Letter, is perhaps the most relevant such model Kevrekidis (2009). It is used to study many physical phenomena involving propagation of nonlinear waves in discrete media, from ultracold gases Trombettoni and Smerzi (2001); Livi et al. (2006); Hennig et al. (2010) in optical lattices to arrays of optical waveguides Christodoulides and Joseph (1988); Lederer et al. (2008). In the DNLS model, energy and norm (or mass) are conserved. For any given mass density, one can construct states of arbitrary energy, but above a critical energy density (corresponding to a so-called infinite temperature), the typical states involve the “condensation” of a finite fraction of energy on a few sites. In other words, the entropy decreases upon increasing energy: the standard signature of a NT.
The interplay between nonlinear dynamics and condensation makes the statistical mechanical description of macroscopic NT states particularly challenging. In spite of the compelling theoretical arguments predicting localization, DNLS microcanonical simulations performed above the infinite-temperature line have shown the emergence of “stationary” spatially homogeneous states Iubini et al. (2013a). Moreover, dynamical studies based on excursion-time distributions Mithun et al. (2018) revealed the occurrence of a transition from ergodic to nonergodic (localized) behavior, strictly above the infinite-temperature line. In the latter regime, localized fast rotations known as discrete breathers induce extremely slow relaxation phenomena Ng et al. (2009); Eckmann and Wayne (2018); Iubini et al. (2019). A partial solution of this paradox was offered in Gradenigo et al. (2021a, b), where it was proven that finite-size corrections are so strong that condensation is inhibited in a strip above the infinite-temperature line. The strip width decreases with the chain length $N$ as $N^{-1/3}$, but is large enough to justify some numerical observations of stationary “non-Gibbsian” states.
So far, this scenario has been discussed in isolated systems; the regime in which mass and energy are exchanged with NT reservoirs received much less attention Baldovin et al. (2021). In this Letter, we revisit the problem by introducing an effective grand-canonical measure, which replaces the non-normalizable standard definition Rasmussen et al. (2000): as a result, we identify a novel class of metastable, delocalized NT states. In particular, we find that the condensation of energy on a single site (breather growth) is dynamically enabled only above a mass threshold $c_{max}\sim|T|$. A properly truncated distribution adequately describes the metastable regime and predicts exponentially long (with $|T|$) lifetimes. The predictions are confirmed by numerical simulations performed with suitable heat baths (after including some precautions to get rid of non-normalizable probabilities Aghion et al. (2019, 2020)). This mechanism, strengthened by the dynamical slowing down of breather states Iubini et al. (2019), clarifies that condensation instabilities can be de facto eluded, making NT states physically accessible also in open setups Kurnosov et al. (2024); Baudin et al. (2023); Iubini et al. (2017a).
Grand-canonical description
In one dimension, the DNLS equation writes
$i\dot{z}_{n}=-2|z_{n}|^{2}z_{n}-z_{n+1}-z_{n-1}\;,$ (1)
where $z_{n}$ is a complex variable, $n=1,\dots,N$ is the index of the lattice site and suitable boundary conditions are assumed (see below). The model has two conserved quantities: the total mass $A=\sum_{n}|z_{n}|^{2}$ and the total energy
$H=\sum_{n=1}^{N}\left(|z_{n}|^{4}+z_{n}^{*}z_{n+1}+z_{n}z_{n+1}^{*}\right),$ (2)
The equilibrium behavior of the model can be described in terms of the mass density $a=A/N$ and the energy density $h=H/N$. Equivalently, the inverse temperature $\beta$ and chemical potential $\mu$ Rasmussen et al. (2000) can be used within the “grand-canonical” representation. Below the infinite-temperature line $h_{c}=2a^{2}$, homogeneous positive-temperature states are found, well described by the Gibbs probability distribution $P_{G}(z_{n})\sim\exp(-\beta H+\beta\mu A)$. Above this line (i.e. for $\beta<0$), the grand-canonical picture is formally ill-defined due to the divergence of $P_{G}(z_{n})$ for arbitrarily large $|z_{n}|$.
Simplified model
Close to $\beta=0$, the hopping energy $H_{int}=\sum_{n}(z_{n}z_{n+1}^{*}+z_{n}^{*}z_{n+1})$ can be neglected in Eq. (2Arezzo et al. (2022). The resulting model, called “C2C” Iubini et al. (2014); Szavits-Nossan et al. (2014); Gradenigo et al. (2021a); Gotti et al. (2021); Arezzo et al. (2022); Gotti et al. (2022), displays a phase-diagram very similar to that of the DNLS equation. It can be implemented as a stochastic evolution involving real variables $c_{n}=|z_{n}|^{2}\geq 0$. A triple of consecutive sites is randomly selected, and the corresponding variables $(c_{n-1},c_{n},c_{n+1})$ are updated in such a way that their sum and the sum of their squares are both conserved. The admissible configurations lie along a circle and the selection is randomly made by assuming a uniform angle distribution. Depending on the initial configuration, the “legal” states (with positive masses) may be restricted to three distinct arcs: in this case, the new state is selected within the same arc of the initial one. Detailed balance is satisfied.
In the positive-temperature region, the invariant measure is the product of single-particle distributions
$P(c)\propto\exp[-(\beta c^{2}-mc)]\quad,$ (3)
[…]
$Z_{r}=-\sqrt{|T|}\tilde{D}\left(m\sqrt{|T|}/2\right)\,,$ (4)
[…]
The average densities $a=\langle A\rangle/N$ and $h=\langle H\rangle/N$ can be thereby derived from the standard grand-canonical relations, $a=\partial_{m}\log(Z_{r})$, $h=\partial_{T}\log(Z_{r})/T^{2}$, obtaining sup
$\begin{cases}a&=\sqrt{|T|}\tilde{D}^{-1}\left(m\sqrt{|T|}/2\right)/2-m|T|/2\\ h&=-(ma+1)|T|/2\,.\end{cases}$ (5)
For $|T|\gg 1$, these expressions reduce to
$\begin{cases}a&=-\frac{1}{m}+\frac{4\beta}{m^{3}}-\frac{40\beta^{2}}{m^{5}}+o(\beta^{2})\\ h&=\frac{2}{m^{2}}-\frac{20\beta}{m^{4}}+o(\beta)\end{cases}$ (6)
which hold independently of the sign of $\beta$. Hence, as long as $c<c_{max}$, the stationary properties change smoothly while crossing the infinite-temperature line. This result is valid also in the thermodynamic limit, but is limited to finite observation times which depend on the (size-dependent) escape time from the interval $[0,c_{max}]$. We now discuss the metastability of this regime.
“Stationary” dynamics
If we formally view $P(c)$ as the outcome of a Langevin process, we can interpret $U(c)\simeq-\ln P(c)=\beta c^{2}-mc$ as an effective potential (complemented by the condition $U(0)=+\infty$). A few samples are presented in Fig. 1: $U(c)$ is binding for $\beta>0$ (see the black curve). For $\beta=0$, $U(c)$ is still binding provided that $m<0$: see the green straight curve, which corresponds to a pure exponential distribution. For $\beta<0$, $U(c)$ is no longer binding, but it is confining for $c<c_{max}$ (the red curve in Fig. 1 corresponds to $\beta=-0.1$).
Figure 1: Effective potential $U(c)$ when passing from $\beta>0$ to $\beta<0$ for fixed $m=-1$.
So far, the interpretation of $U(c)$ is formal. In order to validate its meaning, we consider the restriction of the C2C model to a single triplet $(c_{-},c,c_{+})$, where $c_{-}$ and $c_{+}$ are independent random variables, both distributed according to Eq. (3). It is straightforward to conclude that also the distribution of $c$ is described by Eq. (3) for the same $\beta$ and $m$ values. For NT, $P(c)$ is formally unbounded. However, consistently with our effective grand-canonical approach, it makes sense, slightly above the $\beta=0$ line, to cut the range of allowed $c$ values at $c_{max}$ (i.e. at a minimum of $P(c)$).
Next, we approximate this stochastic system with a nonlinear Langevin process,
$\dot{c}=F(c)+\sqrt{2D(c)}\xi\;,$ (7)
where the effective deterministic drift $F(c)$ and the diffusion coefficient $D(c)$ ($\xi$ is a white noise with unit variance) are empirically defined from numerical simulations
$F(c)=\langle c^{\prime}\rangle-c\quad,\quad 2D(c)=\langle(c^{\prime}-c)^{2}\rangle\;.$ (8)
Here $c^{\prime}$ is the iterate of $c$ as obtained by implementing the C2C model, while the angular brackets denote an average over different realizations of $c_{-}$, $c_{+}$ and of the angle selected in the implementation of the C2C rule.
Figure 2: Velocity field $F(c)$ and diffusion $D(c)$ obtained from numerical simulations of the C2C model for $\beta=0$ (a) and $\beta=-0.01$ (b). In both cases, the corresponding ratio $F(c)/D(c)$ is compared with the expected force $-U^{\prime}(c)$.
The results for infinite temperature and $T=-100$ are plotted in Fig. 2. In both cases $F$ is negative for small $c$ values, indicating that small breathers are stable. A better description is obtained by looking at $F/D$ Gardiner (1985); first of all, we can appreciate that for $\beta=0$, the drift $F(c)$ is negative for all $c$ values (see panel (a)) in agreement with the expectation that breathers of any amplitude are eventually absorbed. For $T=-100$, the drift crosses zero at $c=50$, becoming positive (see the green line in pnale (b)) in agreement with the theoretical expectation $c_{max}=mT/2$ ($m=-1$). On a yet more quantitative level, the asymptotic linear growth of $F/D$ matches the linear dependence of $-U^{\prime}(c)=2\beta c-m$, thus showing that the Langevin equation reproduces the behavior of the C2C model for large $c$ values. In particular, only breathers with $c>c_{max}$ tend to explode, or, equivalently, represent seeds of condensation.
Now, we focus on the lifetime of the metastable states, starting from their stationarity. In Fig. 3 we present the distribution of $c$ values obtained by averaging over 1000 different trajectories until they reach the maximum of the potential at $c_{max}$ (the initial condition having been fixed in $c=1$).
Figure 3: Numerical histograms (solid lines) of $c$ values for different negative $T$ for the C2C model. Dashed lines refer to the corresponding truncated grand-canonical distributions. Upper inset: same analysis for the DNLS chain, Eq. (9). Black curve refers to bulk sites, the red one to the thermalized site. Lower inset: first passage time from the saddle vs $T$ for C2C (full circles) and DNLS (full squares). Dashed lines show the theoretical expectations.
The curves refer to large NT; in all cases the numerical histograms closely follow the theoretical expectation (i.e. the truncated form of Eq. (3)), confirming the validity of the effective grand-canonical approach.
Thanks to the correspondence with the Langevin model, the lifetimes of the metastable state can be estimated via the Kramers formula Risken (1989) for the escape through a barrier located in $c_{max}$, $\tau\approx\exp(\Delta U)/\sqrt{D_{0}D_{1}}$, where $\Delta U$ is the barrier height, while $D_{0}\equiv D(0)$ and $D_{1}\equiv D(c_{max})$ are the diffusion coefficients in the bottom of the valley and at the saddle, respectively. Since $\Delta U=m^{2}|T|/4$, $\tau\approx\exp(\lambda|T|)$, with $\lambda=m^{2}/4$. Numerical simulations performed for $m=-1$ and $T\in[-50,-25]$ yield $\lambda\approx 0.365$ to be compared with the theoretical prediction 0.25. Large part of the discrepancy can be attributed to the indirect dependence of $D_{1}$ on $T$. In fact, as visible in both panels of Fig. 2, $D$ decreases with $c$, while $c_{max}=-T/2$. The effective dependence follows a power law Iubini et al. (2014, 2017b) and thus it does not affect the asymptotic scaling behavior; however, in the range of temperatures here explored, $D_{1}$ contributes an increase of $\lambda$ by about 0.07, bringing the theoretical prediction closer to the observed value (0.32 vs 0.365). The remaining discrepancy is presumably due to the imperfect description of the stochastic dynamics in terms of a continuous nonlinear Langevin process.
DNLS dynamics
Given the above mentioned continuity with the high-temperature region, it makes sense to extend the Langevin scheme proposed in Iubini et al. (2013b) to model the interaction of a DNLS chain with NT reservoirs
$\dot{z}_{n}\!=\!(\gamma_{n}-i)(-2|z_{n}|^{2}z_{n}\!-\!z_{n+1}\!-\!z_{n-1})+\gamma_{n}\mu z_{n}+\sqrt{\gamma_{n}T}\xi_{n}\,$ (9)
where $\gamma_{n}$ is the coupling strength and $\xi_{n}$ a complex Gaussian noise with zero mean and unit variance. While $\gamma_{n}$ is positive for $T>0$, it changes sign at NT. This property is in accord with the general symmetries expected for NT reservoirs Baldovin et al. (2021); $\gamma_{n}<0$ indicates that the sign of nonconservative forces is reversed with respect to the $\beta>0$ case. If we neglect nearest-neighbor interactions in Eq. (9), the single-site non-conservative force $\tilde{F}(z)\simeq\gamma_{n}z_{n}\left(\mu-2|z_{n}|^{2}\right)$ is destabilizing for $|z_{n}|^{2}\geq c_{max}$, as predicted by the single-particle probability $P(c)$, Eq. (3). Similar results had been obtained while studying the stability of breathers in interaction with a bath of linear waves Rumpf (2007, 2009).
In the following we assume periodic boundary conditions ($z_{N+1}=z_{1}$) and thermalize a single site, $\gamma_{n}=\gamma\delta_{n,1}$. The inset of Fig. 4 shows the averages of $a$ and $h$ in a thermostatted DNLS chain with 20 lattice sites, for $T=-100$, $m=-1$ and $\gamma=-0.02$ (in this regime $\tau>10^{9}$ is much longer than the integration time). They are in good quantitative agreement with the truncated grand-canonical prediction Eq. (5) (dashed lines), as also confirmed by the direct extraction of mass distributions, see the upper inset of Fig. 3. This is not a priori obvious, since the theory neglects $H_{int}$. In the metastable $\beta<0$ region of interest for this work $H_{int}$ is indeed negligible sup .
DNLS (frozen) instability
We now study the evolution of discrete breathers determining their drift and diffusion properties as a function of their amplitude $|z^{2}|$. The drift $F(c)$ is estimated by monitoring the evolution of a breather of initial mass $|z|^{2}=c$ in a short ($N=6$) chain thermostatted at a given NT in the antipodal position with respect to the breather site $n_{0}$. To improve numerical accuracy Politi et al. (2022), we monitor the square root of breather energy $\tilde{c}(t)=\{\,|z_{n_{0}}|^{4}+[z_{n_{0}}^{*}(z_{n_{0}+1}+z_{n_{0}-1})+c.c.]\,\}^{1/2}$, (dimensionally equivalent to the local mass), and compute $F(c)=\langle d\tilde{c}/dt\rangle$. Analogously, $D(c)=Var[\,\tilde{c}(t)\,]/2t$, where $Var(\cdot)$ denotes the variance of the signal. The red crosses reported in Fig. 4 show that the drift is negative for small $c$-values, indicating that such breathers are stable, consistently with the numerical evidence of pseudo stationary regimes. However, at variance with the stochastic C2C model, here $F(c)$ becomes indistinguishable from zero above $c\approx 17$. This slowing down is a manifestation of the adiabatic invariant (AI) unveiled in Iubini et al. (2019); Politi et al. (2022) in the positive-temperature region.
In order to bypass this obstacle, we have also analyzed a stochastic version of the DNLS equation (SDNLS), where the AI is absent. It amounts to including the sporadic update of the phase of a randomly chosen $z_{n}$ so as to leave the local interaction energy unchanged sup ; Iubini (2019). The SDNLS exhibits the same stationary properties as the DNLS model, but is characterized by a substantially faster breather dynamics.
Figure 4: (a,b) Relaxation to pseudo-equilibrium of a DNLS equation interacting with a generalized Langevin reservoir with $\beta=-0.01$, $m=-1$ and $\gamma=-0.02$: evolution of average mass $a(t)$ and average energy $h(t)$. Red dashed lines refer to the analytical prediction in (5). (c) Effective force $F$ (scaled by $c$ for clarity) acting on a breather with initial mass $c$ for $m=-1$ and different $T$. Open symbols refer to the SDNLS model; red crosses to the full DNLS for $T=-34$. (d) Ratio $F/D$ compared to theoretical values of $-U^{\prime}(c)$ (dashed lines). Same symbols as in (c). Inset: DNLS evolution of $\tilde{c}(t)$ for $c=21$. The dot-dashed line shows the average trend excluding downward bursts.
Simulations of the SDNLS chain for different temperatures show that the drift is now significantly different from 0 and quantitatively measurable (see circles and triangles in Fig. 4(c)). In particular, one can identify a crossing point above which the drift is positive. Also the diffusion coefficient can be estimated; the resulting numerical values of $F/D$ are presented in Fig. 4(d). They are relatively close to the theoretical predictions (see the dashed lines); the main difference is a rightward shift which implies that the numerically estimated saddle point $\tilde{c}_{max}$ is slightly larger than the expected value (on the basis of the C2C model). Anyway, the deviation $\tilde{c}_{max}-c_{max}$ does not increase upon increasing $|T|$ sup .
A fully quantitative analysis of the deterministic DNLS equation is not doable. However, we have made our best efforts to determine $F/D$ in a single point in the presence of an AI. We have performed a single long numerical simulation for $T=-34$ and an initial breather mass $c=21>c_{max}=17$ $N=12$: the evolution of $\tilde{c}(t)$ is reported in the inset of Fig. 4(d). It reveals large bursts and less visible but relevant blobs which all hinder a quantitative analysis, since these localized instabilities do not typically contribute to the overall growth and diffusion  [^1] They signal the emergence of “dimer” bound states when the energy temporarily spreads over two adjacent sites and yet strongly affect numerical estimates. After a lengthy manual removal of the most relevant bursts, we have obtained a cleaner curve  sup , whose analysis leads to $F/D\approx 0.38$, in semi-quantitative agreement with the theoretical expectation (see the uppermost red cross in Fig. 4(d)).
Finally, we consider the escape time out of the DNLS metastable regime. The major difference with the C2C model is the dependence of the diffusion coefficient $D$ on the mass contained in the breather. As shown in Iubini et al. (2019); Politi et al. (2022), the AI induces an exponential decrease, $D\approx\exp(-\alpha c)$. Hence, $D_{1}\simeq\exp(-\alpha mT/2)$, which, inserted into the Kramers formula, enhances the exponential growth rate of the lifetime, to $\lambda\approx m^{2}/4+\alpha|m|/4$. This implies (for $m=-1$), $\lambda\approx 0.5$. Direct numerical simulations performed for $N=7$ yield the results reported in Fig. 3, an exponential fit gives a rate 0.53 in good agreement with the theoretical expectation.
Conclusions
Recent large-deviation studies, based on a microcanonical analysis of finite systems at infinite times, pointed out the existence of a first-order condensation transition at $\beta=0$, accompanied by relevant finite-$N$ corrections on the NT side. Here, we have revisited the problem by implementing a grand-canonical approach, thanks to the introduction of thermal reservoirs, which operate at large NTs. Our key result is the evidence of a temperature-dependent critical amplitude, separating stable from unstable condensation peaks (DNLS breathers). As a result, close to the $\beta=0$ line, extended states are metastable but characterized by an exponentially long lifetime. In the DNLS model, the exponential growth is enhanced by the fast decrease of the diffusion coefficient with the breather height. In this Letter we have assumed the presence of a single breather; in general, multiple breathers are expected to arise in the NT region, each accompanied by an additional quasi-conserved quantity (the AI). It is conceivable that a generalized Gibbs ensembles should be implemented, analogous to what done in the context of integrable systems Vidmar and Rigol (2016); Spohn (2020). This generalization is, however, not straightforward; it requires further work.
Acknowledgements.
We thank P. Politi for fruitful discussions.