### v2

# Effective grand-canonical description of condensation in negative-temperature regimes

Abstract
The observation of negative-temperature states in the localized phase of the Discrete Nonlinear Schrödinger (DNLS) equation has challenged statistical mechanics for a long time. For isolated systems, they can emerge as stationary extended states through a large-deviation mechanism occurring for finite sizes, while they are formally unstable in grand-canonical setups, being associated to an unlimited growth of the condensed fraction. Here, we show that negative-temperature states in open setups are metastable and their lifetime $\tau$ is exponentially long with the temperature, $\tau\approx\exp(\lambda|T|)$ (for $T<0$). A general expression for $\lambda$ is obtained in the case of a simplified stochastic model of non-interacting particles. In the DNLS model, the presence of an adiabatic invariant, makes $\lambda$ even larger because of the resulting freezing of the breather dynamics. This mechanism, based on the existence of two conservation laws, provides a new perspective over the statistical description of condensation processes.
Statistical systems exhibiting negative (absolute) temperatures (NTs) often display awkward physical properties, but a careful analysis shows that their behavior does not contradict any fundamental law of thermodynamics Baldovin et al. (2021). NTs typically emerge when the total energy is somehow limited from above Ramsey (1956), but they can occur also in systems with unbounded energies, provided that the first moment of the density of states in energy remains finite Machlup (1975). Lasers are a clear example of the first kind: maximal energy is attained only by setting all the atoms in the excited state. Vortices in bounded fluids are an example of the second kind Onsager (1949); Gauthier et al. (2019); Johnstone et al. (2019). Other recent observations of NT states include cold atoms in optical lattices Braun et al. (2013) and multimode optical fibers Baudin et al. (2023); Marques Muniz et al. (2023).
Energy limitation can also arise indirectly in the presence of additional conservation laws. The discrete nonlinear Schrödinger (DNLS) equation, the subject of this Letter, is perhaps the most relevant such model Kevrekidis (2009). It is used to study many physical phenomena involving propagation of nonlinear waves in discrete media, from ultracold gases Trombettoni and Smerzi (2001); Livi et al. (2006); Hennig et al. (2010) in optical lattices to arrays of optical waveguides Christodoulides and Joseph (1988); Lederer et al. (2008). In the DNLS model, energy and norm (or mass) are conserved. For any mass density, one can construct states of arbitrary energy, but above a critical energy density (corresponding to a so-called infinite temperature), the entropy decreases upon increasing energy: the standard signature of a NT. As a result, typical states involve the “condensation” of a finite fraction of energy on a few sites.
Nevertheless, in Gradenigo et al. (2021a, b), it was proven that finite-size corrections are so strong that condensation is inhibited in a strip of width $N^{-1/3}$ ($N$ being the chain length), above the infinite-temperature line. The strip is large enough to justify some numerical results, such as the observation of “stationary” spatially homogeneous states Iubini et al. (2013a) and of “non-Gibbsian” states Mithun et al. (2018) (the latter ones detected via a careful analysis of excursion-time distributions).
The complexity of the underlying regime can be traced back to the emergence of localized fast rotations known as discrete breathers, around which a sensible amount of energy can condense. In the past, breathers have been studied as single dynamical objects, by investigating the stability of energy clumps Trombettoni and Smerzi (2001); Hennig et al. (2016); Franzosi et al. (2011). Here, we aim at characterizing breather dynamics in a statistical-mechanics setup, by singling out the role of thermodynamic observables.
A statistical description of localization stability has been already produced for breathers interacting with a bath of waves in terms of a matching condition between breather frequency and the chemical potential of the wave background Rumpf (2007, 2009); Chen and Rumpf (2021). Furthermore, tall (i.e. rapidly rotating) breathers have been found to undergo extremely slow relaxation phenomena Ng et al. (2009); Eckmann and Wayne (2018); Iubini et al. (2019).
So far, this problem was discussed in isolated systems. In this Letter, we study a setup where mass and energy are exchanged with NT reservoirs, by introducing an effective grand-canonical measure, which replaces the non-normalizable standard definition Rasmussen et al. (2000). As a result, we identify a novel class of extended NT states and describe the stationary regimes in terms of a properly truncated distribution. With the help of an effective Langevin equation (after including some precautions to get rid of non-normalizable probabilities Aghion et al. (2019, 2020)), the steady description is inserted into a dynamical context, which leads to two predictions: (i) condensation progresses (on average) only above a critical mass threshold $c_{max}\sim|T|$, in agreement with the results in Rumpf (2007, 2009); Chen and Rumpf (2021); (ii) the lifetime of the associated metastable state is exponentially long (with $|T|$). The lifetime is lenghtened by the dynamical slowing down of tall breathers Iubini et al. (2019), de facto eluding the condensation instabilities and making NT states physically accessible also in open setups Kurnosov et al. (2024); Baudin et al. (2023); Iubini et al. (2017a).
Grand-canonical description
In one dimension, the DNLS equation writes
$i\dot{z}_{n}=-2|z_{n}|^{2}z_{n}-z_{n+1}-z_{n-1}\;,$ (1)
where $z_{n}$ is a complex variable, $n=1,\dots,N$ is the index of the lattice site and suitable boundary conditions are assumed (see below). The model has two conserved quantities: the total mass $A=\sum_{n}|z_{n}|^{2}$ and the total energy  [^1] Sometime in the literature, the quartic term is preceded by a coefficient $\nu$ quantifying the nonlinearity Rasmussen et al. (2000). It can be eliminated via a suitable rescaling. Its explicit presence implies that the expression of the critical threshold $c_{max}$ should be divided by $\nu$ (see below).
$H=\sum_{n=1}^{N}\left(|z_{n}|^{4}+z_{n}^{*}z_{n+1}+z_{n}z_{n+1}^{*}\right),$ (2)
The equilibrium behavior of the model can be described in terms of the mass density $a=A/N$ and the energy density $h=H/N$. Equivalently, the inverse temperature $\beta$ and chemical potential $\mu$ Rasmussen et al. (2000) can be used within the “grand-canonical” representation. Below the infinite-temperature line $h_{c}=2a^{2}$, homogeneous positive-temperature states exista and are well described by the Gibbs probability distribution $P_{G}(z_{n})\sim\exp(-\beta H+\beta\mu A)$. Above this line (i.e. for $\beta<0$), the grand-canonical picture is formally ill-defined due to the divergence of $P_{G}(z_{n})$ for arbitrarily large $|z_{n}|$.
Simplified model
Close to $\beta=0$, the hopping energy $H_{int}=\sum_{n}(z_{n}z_{n+1}^{*}+z_{n}^{*}z_{n+1})$ can be neglected in Eq. (2Arezzo et al. (2022). The resulting model, called “C2C” Iubini et al. (2014); Szavits-Nossan et al. (2014a); Gradenigo et al. (2021a); Gotti et al. (2021); Arezzo et al. (2022); Gotti et al. (2022), displays a phase-diagram very similar to that of the DNLS equation. It can be implemented as a stochastic evolution involving real variables $c_{n}=|z_{n}|^{2}\geq 0$ in such a way that their sum and the sum of their squares are both conserved sup .
In the positive-temperature region, the invariant measure is the product of single-particle distributions
$P(c)\propto\exp[-(\beta c^{2}-mc)]\quad,$ (3)
[…]
$Z_{r}=-\sqrt{|T|}\tilde{D}\left(m\sqrt{|T|}/2\right)\,,$ (4)
[…]
The average densities $a=\langle A\rangle/N$ and $h=\langle H\rangle/N$ can be derived from the grand-canonical relations, $a=\partial_{m}\log(Z_{r})$, $h=\partial_{T}\log(Z_{r})/T^{2}$ (see sup ). For $|T|\gg 1$, the expressions reduce to
$\begin{cases}a&=-\frac{1}{m}+\frac{4\beta}{m^{3}}-\frac{40\beta^{2}}{m^{5}}+o(\beta^{2})\\ h&=\frac{2}{m^{2}}-\frac{20\beta}{m^{4}}+o(\beta)\end{cases}$ (5)
which hold independently of the sign of $\beta$. Hence, as long as $c<c_{max}$, the stationary properties (including those of spatial profiles) change smoothly while crossing the infinite-temperature line.
One can formally interpret $U(c)\simeq-\ln P(c)=\beta c^{2}-mc$ as the effective potential (complemented by the condition $U(0)=+\infty$) in an underlying Langevin process. A few instances are presented in Fig. 1: $U(c)$ is binding for $\beta>0$ (see the black curve). For $\beta=0$, $U(c)$ is still binding provided that $m<0$: see the green straight curve, which corresponds to a pure exponential distribution. For $\beta<0$, $U(c)$ is no longer binding, but it is confining for $c<c_{max}$ (the red curve in Fig. 1 corresponds to $\beta=-0.1$).
Figure 1: Effective potential $U(c)$ when passing from $\beta>0$ to $\beta<0$ for fixed $m=-1$.
In order to validate this formal interpretation, we consider the restriction of the C2C model to a single triplet $(c_{-},c,c_{+})$, where $c_{-}$ and $c_{+}$ are independent random variables, both distributed according to Eq. (3). It is straightforward to conclude that also the distribution of $c$ is described by Eq. (3) for the same $\beta$ and $m$ values. For NT, $P(c)$ is formally unbounded. However, consistently with our effective grand-canonical approach, it makes sense, slightly above the $\beta=0$ line, to cut the range of allowed $c$ values at $c_{max}$ (i.e. at a minimum of $P(c)$).
In Fig. 2 we present the distribution of $c$ values obtained by averaging over 1000 different trajectories sampled until they reach $c_{max}$ (the initial condition having been fixed in $c=1$).
Figure 2: Numerical histograms (solid lines) of $c$ values for different negative $T$ for the C2C model. Dashed lines refer to the corresponding truncated grand-canonical distributions. Upper inset: same analysis for the DNLS chain, Eq. (8). Black curve refers to bulk sites, the red one to the thermalized site. Lower inset: first passage time from the saddle vs $T$ for C2C (full circles) and DNLS (full squares). Dashed lines show the theoretical expectations.
The curves refer to large NT; in all cases the numerical histograms closely follow the theoretical expectation (i.e. the truncated form of Eq. (3)), confirming the validity of the effective grand-canonical approach.
Effective Langevin description
Here, we approximate the model dynamics with the stochastic equation
$\dot{c}=F(c)+\sqrt{2D(c)}\xi\;,$ (6)
where the deterministic drift $F(c)$ and the diffusion coefficient $D(c)$ ($\xi$ is a white noise with unit variance) are empirically defined from numerical simulations
$F(c)=\langle c^{\prime}\rangle-c\quad,\quad 2D(c)=\langle(c^{\prime}-c)^{2}\rangle\;.$ (7)
Here $c^{\prime}$ is the iterate of $c$ as obtained by implementing the C2C model. The angular brackets denote an average over all possible realizations of the thermal baths and implementations of the C2C rule.
Figure 3: Velocity field $F(c)$ and diffusion $D(c)$ obtained from numerical simulations of the C2C model for $\beta=0$ (a) and $\beta=-0.01$ (b). In both cases, the corresponding ratio $F(c)/D(c)$ is compared with the expected force $-U^{\prime}(c)$.
In Fig. 3, we see that both for $\beta=0$ and $-0.01$, $F$ is negative for (not too) small $c$ values, indicating that small breathers are stable. $F/D$ Gardiner (1985) provides a better representation. For $\beta=0$, we see that the drift is still negative at large $c$ values (see panel (a)) in agreement with the expectation that breathers are eventually absorbed. For $\beta=-0.01$, the drift crosses zero at $c=50$, becoming positive (see the green line in panel (b)) in agreement with the theoretical expectation $c_{max}=mT/2$ ($m=-1$). On a yet more quantitative level, the asymptotic linear growth of $F/D$ matches the linear dependence of $-U^{\prime}(c)=2\beta c-m$, thus showing that the Langevin equation reproduces the behavior of the C2C model for large $c$ values. In particular, only breathers with $c>c_{max}$ tend to explode, or, equivalently, represent seeds of condensation.
Once $F$ and $D$ have been determined, one can estimate the lifetime of the metastable state via Kramers formula Risken (1989) for the escape through a barrier located in $c_{max}$, $\tau\approx\exp(\Delta U)/\sqrt{D_{0}D_{1}}$, where $\Delta U$ is the barrier height, while $D_{0}\equiv D(0)$ and $D_{1}\equiv D(c_{max})$ are the diffusion coefficients in the bottom of the valley and at the saddle, respectively. Since $\Delta U=m^{2}|T|/4$, $\tau\approx\exp(\lambda|T|)$, with $\lambda=m^{2}/4$. Numerical simulations performed for $m=-1$ and $T\in[-50,-25]$ yield $\lambda\approx 0.365$ to be compared with the theoretical prediction 0.25. Large part of the discrepancy can be attributed to the indirect dependence of $D_{1}$ on $T$. In fact, as visible in both panels of Fig. 3, $D$ decreases with $c$, while $c_{max}=-T/2$. The effective dependence follows a power law Iubini et al. (2014, 2017b) and thus it does not affect the asymptotic scaling behavior; however, in the range of temperatures here explored, $D_{1}$ contributes an increase of $\lambda$ by about 0.07, bringing the theoretical prediction closer to the observed value (0.32 vs 0.365). The remaining discrepancy is presumably due to the imperfect description of the stochastic dynamics in terms of a continuous nonlinear Langevin process.
DNLS dynamics
Given the above mentioned continuity with the high-temperature region, it makes sense to extend the scheme proposed in Iubini et al. (2013b) to model the interaction of a DNLS chain with NT reservoirs
$\dot{z}_{n}\!=\!(\gamma_{n}-i)(-2|z_{n}|^{2}z_{n}\!-\!z_{n+1}\!-\!z_{n-1})+\gamma_{n}\mu z_{n}+\sqrt{\gamma_{n}T}\xi\,$ (8)
where $\gamma_{n}=\delta_{n1}\gamma$ is the coupling strength and $\xi$ a complex Gaussian noise with zero mean and unit variance. Periodic boundary conditions ($z_{N+1}=z_{1}$) are assumed. While $\gamma$ is positive for $T>0$, it changes sign at NT. This property is in accord with the general symmetries expected for NT reservoirs Baldovin et al. (2021); $\gamma<0$ indicates that the sign of nonconservative forces is reversed with respect to the $\beta>0$ case. If we neglect the nearest-neighbor interactions of the thermostatted site, the evolution equation of the mass $c=|z_{1}|^{2}$ writes $\dot{c}=-2\gamma(2c^{2}-\mu)c+z^{*}\xi+z\xi^{*}$, a form similar to the Langevin Eq. (6), with the difference that this equation follows from a precise thermal-bath definition Iubini et al. (2013b) instead of being a mere conjecture. It is also interesting to notice that the effective force $\tilde{F}(c)=2\gamma c\left(\mu-2c^{2}\right)$ predicts a destabilization for $c^{2}\geq c_{max}$, consistent with the single-particle probability $P(c)$, Eq. (3).
Fig. 4 (a,b) shows the averages of $a$ and $h$ in a DNLS chain of length $N=20$, thermostatted for $T=-100$, $m=-1$ and $\gamma=-0.02$ (in this regime $\tau>10^{9}$ is much longer than the integration time). They are in good quantitative agreement with the truncated grand-canonical prediction Eq. (5) (dashed lines), as also confirmed by the direct extraction of mass distributions, see the upper inset of Fig. 2. This correspondence is not a priori obvious, since the theory neglects $H_{int}$. In the metastable $\beta<0$ region of interest for this work, $H_{int}$ is on average negligible sup .
DNLS (frozen) instability
We now study the evolution of discrete breathers for different amplitudes $|z^{2}|$. At variance with the C2C model, here thermostats induce strong boundary effects. As previously shown in positive-temperature simulations Politi et al. (2022), the high frequencies naturally generated by the heat bath spuriously affect the DNLS dynamics (in particular the breather stability). Hence, Eq. (8) cannot account for the true dynamics of the background. Nevertheless, Eq. (6) provides again a good effective description. By determining the drift $F$ and diffusion $D$ of breathers of mass $|z|^{2}=c$ (located in the antipodal position with respect to the thermal bath) for different chain lengths, we have verified that a length $N=6$ suffices to kill unphysical high frequencies.
To improve numerical accuracy Politi et al. (2022), we monitor the square root of breather energy $\tilde{c}(t)=\{\,|z_{n_{0}}|^{4}+[z_{n_{0}}^{*}(z_{n_{0}+1}+z_{n_{0}-1})+c.c.]\,\}^{1/2}$, (dimensionally equivalent to the local mass), and compute $F(c)=\langle d\tilde{c}/dt\rangle$. Analogously, $D(c)=Var[\,\tilde{c}(t)\,]/2t$, where $Var(\cdot)$ denotes the variance of the signal. The red crosses reported in Fig. 4 show that the drift is negative for small $c$-values, indicating that such breathers are stable, consistently with the numerical evidence of pseudo stationary regimes. However, at variance with the stochastic C2C model, here $F(c)$ becomes indistinguishable from zero above $c\approx 17$. This slowing down is a manifestation of the adiabatic invariant unveiled in Iubini et al. (2019); Politi et al. (2022) in the positive-temperature region.
In order to bypass this obstacle, we have also analyzed a stochastic version of the DNLS equation (SDNLS), where the adiabatic invariant is absent. It amounts to including the sporadic update of the phase of a randomly chosen $z_{n}$ so as to leave the local interaction energy unchanged sup ; Iubini (2019). The SDNLS exhibits the same stationary properties as the DNLS model, but is characterized by a substantially faster breather dynamics.
Figure 4: (a,b) Relaxation to pseudo-equilibrium of a DNLS equation interacting with a generalized Langevin reservoir with $\beta=-0.01$, $m=-1$ and $\gamma=-0.02$: evolution of average mass $a(t)$ and average energy $h(t)$. Red dashed lines refer to the analytical prediction in (5). (c) Effective force $F$ (scaled by $c$ for clarity) acting on a breather with initial mass $c$ for $m=-1$ and different $T$. Open symbols refer to the SDNLS model; red crosses to the full DNLS for $T=-34$. (d) Ratio $F/D$ compared to theoretical values of $-U^{\prime}(c)$ (dashed lines). Same symbols as in (c). Inset: DNLS evolution of $\tilde{c}(t)$ for $c=21$. The dot-dashed line shows the average trend excluding downward bursts.
Simulations of the SDNLS chain show that the drift is now significantly different from 0 and quantitatively measurable (see circles and triangles in Fig. 4(c)). In particular, one can identify a crossing point above which the drift is positive. Also the diffusion coefficient can be estimated; the resulting numerical values of $F/D$ are presented in Fig. 4(d). They are relatively close to the theoretical predictions (see the dashed lines); the main difference is a rightward shift which implies that the numerically estimated saddle point $\tilde{c}_{max}$ is slightly larger than the expected value (on the basis of the C2C model). Anyway, the deviation $\tilde{c}_{max}-c_{max}$ does not increase upon increasing $|T|$ sup .
A fully quantitative analysis of the deterministic DNLS equation is not doable. However, we have made our best efforts to determine $F/D$ in a single point in the presence of an adiabatic invariant. We have performed a single long numerical simulation for $T=-34$ and an initial breather mass $c=21>c_{max}=17$ for $N=12$: the evolution of $\tilde{c}(t)$ is reported in the inset of Fig. 4(d). It reveals large bursts and less visible but relevant blobs which all hinder a quantitative analysis, since these localized instabilities do not typically contribute to the overall growth and diffusion  [^2] They signal the emergence of “dimer” bound states when the energy temporarily spreads over two adjacent sites and yet strongly affect numerical estimates. After a lengthy manual removal of the most relevant bursts, we have obtained a cleaner curve  sup , whose analysis leads to $F/D\approx 0.38$, in semi-quantitative agreement with the theoretical expectation (see the uppermost red cross in Fig. 4(d)).
Finally, we consider the escape time out of the DNLS metastable regime. The major difference with the C2C model is the dependence of the diffusion coefficient $D$ on the mass contained in the breather. As shown in Iubini et al. (2019); Politi et al. (2022), the adiabatic invariant induces an exponential decrease, $D\approx\exp(-\alpha c)$, with $\alpha\approx 1$. Hence, $D_{1}\simeq\exp(-\alpha mT/2)$, which, inserted into the Kramers formula, enhances the exponential growth rate of the lifetime, to $\lambda\approx m^{2}/4+\alpha|m|/4$. This implies (for $m=-1$), $\lambda\approx 0.5$. Direct numerical simulations performed for $N=7$ yield the results reported in Fig. 2; an exponential fit gives a rate 0.53 in good agreement with the theoretical expectation.
Conclusions
We have revisited the stability of extended states by implementing a grand-canonical approach and introducing suitable thermal reservoirs, which operate at large NTs. Our main achievement is the demonstration of metastable states characterized by exponentially long lifetimes. This result follows from the existence of a temperature-dependent critical amplitude, which separates stable from unstable condensation peaks (DNLS breathers). In the DNLS model, the exponential growth is amplified by the fast decrease of the diffusion coefficient with the breather height.
In this Letter, we focused on a setup where the local onsite energy $\epsilon$ is a quadratic function of the mass $c$, but the formalism can be applied to contexts where $\epsilon(c)>0$ (see, e.g., purely stochastic condensation models Szavits-Nossan et al. (2014a, b)): in such cases $c_{max}$ solves $\beta\epsilon^{\prime}(c_{max})=m$. Finally, we have assumed the presence of a single breather; in general, multiple breathers are expected to arise in the NT region, each accompanied by an additional quasi-conserved quantity. It is conceivable that a generalized Gibbs ensemble should be implemented, analogous to what done in the context of integrable systems Vidmar and Rigol (2016); Spohn (2020). This generalization is, however, not straightforward; it requires further work.
Acknowledgements.
We thank P. Politi for fruitful discussions. We acknowledge support from the project “Breakdown of ergodicity in classical and quantum many-body systems” (BECQuMB) funded by the Italian Ministry for University and Research (MUR) PRIN2022 grant No. 20222BHC9Z.